# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ 3.2 Traditional Self-play Algorithms
3.2.1 Integration into Our Framework
Traditional self-play algorithms can be incorporated into our proposed framework (Algo. 1) with the following settings. First, the policy population $\Pi$ utilizes lazy initialization because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy. Here, the policy population size $N$ serves as the upper limit for the number of active policies in the population. In other words, we use $N$ iterations to optimize the policy. Third, the strategy being trained $\pi_{i}^{h}$ can be initialized in a general manner. For instance, the strategy can be initialized randomly, learning from scratch. More often, $\pi_{i}^{h}$ is initialized by $\pi_{i-1}(\cdot|h(i-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(i)=\emptyset$. Fifth, the MSSes of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation. Only the MSS for prioritized fictitious self-play (described in Sec. 3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes like NE.
@@ 3.2.1 Integration into Our Framework
In a two-player symmetric game, $C_{1}=N$, with $\sigma_{ij}$ denoting the probability that policy $i$ is optimized in response to policy $j$ which leads to $\sum_{j=1}^{N}\sigma_{ij}=1,~{}\forall i$.
@@
The policy population gradually increases over time. When policy $i$ is selected for training, only policy $j$ $(j\leq i)$ has already been trained and holds meaningful outcomes. Other policies are inactive policies. As a result, we exclusively select policy $j$, where $j\leq i$, to serve as the opponent for policy $i$. ∎
3.2.2 Vanilla Self-play
@@
In vanilla self-play (Samuel, 2000), agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
@@
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}1,~{}\text{if}~{}j=i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (5)
@@
Regardless of $\mathcal{P}$, the MSS produces the same interaction matrix. Although this MSS is straightforward, it is utilized in many other algorithms, so we refer to it as vanilla MSS. Vanilla self-play is effective in transitive games, but it can lead the agent to cyclic learning patterns in non-transitive games. It may also lead to an unstationary problem, causing the system to get stuck in a local optimum.
@@
A more modern version of vanilla self-play is to combine it with off-policy RL. In each iteration, the policy to be trained, $\pi_{i}^{h}$, is initialized using the previous policy $\pi_{i-1}(\cdot|h(i-1))$. Typically, each player’s policy is fixed, and samples are collected into a buffer by vanilla self-play. These samples remain useful for training, even if they were generated by an earlier policy. This allows the agent to leverage past experiences and maintain more stationary training. Once the policy is updated, it will be added to the policy population.
3.2.3 Fictitious Self-play
@@
Fictitious Play (FP) (Brown, 1951) is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-Play (FSP) (Heinrich et al., 2015) is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-Play (NFSP) (Heinrich & Silver, 2016) is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. In the original versions of NFSP and FSP, two distinct types of agent memory are used: one to record the agent’s own behavior and the other to capture the opponent’s behavior.
@@
In more recent approaches (Lanctot et al., 2017), random sampling is frequently employed to approximate the opponents’ average strategy, eliminating the need to maintain two separate types of agent memory. Thus, the MSS of FSP:
@@
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{i-1},~{}\text{if}~{}j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right..$ (6)
@@
In FSP, the MSS continues to generate a constant interaction matrix. Compared to vanilla self-play, this approach enhances the robustness of the policy by sampling older versions of its own policies to be used as opponent strategies.
3.2.4 $\delta$-uniform Self-play
[…]
@@ 3.2.4 $\delta$-uniform Self-play
$\textit{MSS}(\mathcal{P})_{ij}=\left\{\begin{array}[]{l}\frac{1}{f(i)},~{}\text{if}~{}\lceil\delta(i-1)\rceil\leq j\leq i-1\vspace{4pt}\\ 0,~{}\text{otherwise}\end{array}\right.,$ (7)
$f(i)=i-\lceil\delta(i-1)\rceil.$ (8)
@@
In $\delta$-uniform self-play, The MSS generates a constant interaction matrix. Specially, if $\delta=0$, it corresponds to FSP, and if $\delta=1$, it corresponds to vanilla self-play.
3.2.5 Prioritized Fictitious Self-play
@@
Prioritized Fictitious Self-Play (PFSP)  (Vinyals et al., 2019) leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{ij}=\text{Prob}(\pi_{i}~{}\text{beats}~{}\pi_{j})$. The MSS of PFSP is given by Algo. 5. The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, $f=x(1-x)$ implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics. A similar MSS in  Berner et al. (2019) can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
@@
2:   $\sigma_{i+1,j}\leftarrow\frac{f(\mathcal{P}_{i,j})}{\sum_{j\leq i}f(\mathcal{P}_{i,j})},~{}\text{if}~{}j\leq i$
3:   Append zeros to $\sigma_{i+1}$ until its length is N.
@@ 3.3 PSRO Series
3.3.1 Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$. Second, we also set $E=1$ and $N$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{i}^{h}$ can also be initialized in a general manner. Fourth, we set $h(i)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model (Lanctot et al., 2017) in how $\sigma_{i}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{i}$ is the opponent sampling strategy. It means that $\sigma_{i}$ here represents the opponent’s meta-strategy against policy $i$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSes of the PSRO series are often more complex. For example, some MSSes incorporate concepts from different types of game equilibria (McMahan et al., 2003; Muller et al., 2020; Marris et al., 2021). Lastly, we also simply follow Assumption 1. We can derive the Corollary 2 using a similar proof as Corollary 1.
@@ 3.3.1 Integration into Our Framework
3.3.2 Double Oracle
Double Oracle (DO) (McMahan et al., 2003) can be only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{i}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In the context of two-player normal-form games, DO theoretically can achieve the NE of the full game.
@@ 3.3.2 Double Oracle
2:   $\sigma_{i+1}\leftarrow$ SOLVE-NASH$(\mathcal{P}_{1:i,1:i})$triangleright Opponent’s NE meta-strategy.
3:   Append zeros to $\sigma_{i+1}$ until its length is N.
@@
We take rock-paper-scissors as an example, a symmetric two-player game. The policy condition function $h(i)=\emptyset$. The policy population size $N=4$. The policy population $\Pi$ uses lazy initialization and the interaction matrix $\Sigma$ is initialized with all zeros. Assume that in the first iteration, the policy $\pi_{1}$ is set to "rock". Obviously, the NE-based MSS returns $\sigma_{2}=(1,0,0,0)$. In the second iteration, the BR oracle adds "paper" to $\Pi$ as $\pi_{2}$. Consequently, the NE-based MSS yields $\sigma_{3}=(0,1,0,0)$, since "paper" always beats "rock." In the third iteration, noting that $\sigma_{3}$ selects only "paper" and never "rock," the BR oracle adds "scissors" as $\pi_{3}$. The NE-based MSS then solves the game and returns $\sigma_{4}=(1/3,1/3,1/3,0)$. In the final iteration, the BR oracle adds a mixed strategy that assigns a uniform probability over "rock," "paper," and "scissors" as $\pi_{4}$. The final interaction matrix is:
$\Sigma=\begin{bmatrix}0&0&0&0\\ 1&0&0&0\\ 0&1&0&0\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{3}&0\end{bmatrix}$ (9)
@@
This process demonstrates how the DO algorithm iteratively expands the strategy set within our framework. Also, using DO as an example, we show that, compared with traditional self-play algorithms, the PSRO series employs more sophisticated MSSes to better capture the essence of the game, albeit at a higher computational cost.
3.3.3 PSRO
@@
PSRO (Lanctot et al., 2017) extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSes to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
3.3.4 PSRO Variants
@@
The traditional PSRO algorithm has been the subject of numerous extensions in recent research. Some studies focus on making PSRO more computationally tractable in different scenarios or achieving fast convergence. By establishing a hierarchical structure of RL workers, Pipeline PSRO (McAleer et al., 2020) achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO (Zhou et al., 2022) introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO. Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) (McAleer et al., 2021) allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games. Online Double Oracle (ODO) (Dinh et al., 2022) integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff. Anytime PSRO (McAleer et al., 2022b) and Self-play PSRO (McAleer et al., 2022a) are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence. Moreover, as the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO (Muller et al., 2022) has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games (Daskalakis et al., 2009; Daskalakis, 2013) and the selection problem in NE (Goldberg et al., 2013), Müller et al. put forward $\alpha$-PSRO (Muller et al., 2020). Instead of NE, $\alpha$-rank (Omidshafiei et al., 2019), which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, Marris et al. (2021) proposes Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games. It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
@@
Other research focuses on policy diversification, as deriving a single policy in transitive games is less meaningful. Balduzzi et al. (2019) introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. Perez-Nieves et al. (2021) uses determinantal point process (Kulesza et al., 2012) to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO. Similarly,  Liu et al. (2021) introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO (Yao et al., 2023) defines a diversity metric named population exploitability that helps to achieve a full-game NE.
3.3.5 R-NaD

### v4

@@ 3.2. Traditional Self-Play Algorithms
3.2.1. Integration into Our Framework
Traditional self-play algorithms can be incorporated into our proposed framework (Algo. 1) with the following settings. First, the policy population $\Pi$ utilizes lazy initialization because the policy population in traditional self-play algorithms is intended to grow with each iteration. Second, we set $E=1$ because only the inactive policy can be trained in each iteration, then turning into an active policy. Here, the policy population size $K$ serves as the upper limit for the number of active policies in the population. In other words, we use $K$ iterations to optimize the policy. Third, the strategy being trained $\pi_{[k]}^{h}$ can be initialized in a general manner. For instance, the strategy can be initialized randomly, learning from scratch. More often, $\pi_{[k]}^{h}$ is initialized by $\pi_{[k-1]}(\cdot|h(k-1))$, allowing incremental learning and adaptation based on the most current trained policy to accelerate convergence. Fourth, the policies in traditional self-play algorithms don’t need supplementary information, we set the policy condition function $h(k)=\emptyset$. Fifth, the MSSs of traditional self-play algorithms are straightforward, often yielding a constant interaction matrix $\Sigma$ that eliminates performance evaluation. Only the MSS for prioritized fictitious self-play (described in Sec. 3.2.5) requires the performance matrix; nonetheless, it does not need to solve for complex game outcomes like NE.
@@ 3.2.1. Integration into Our Framework
In a two-player symmetric game, $C_{1}=K$, with $\Sigma_{mn}$ denoting the probability that policy $m$ is optimized in response to policy $n$ which leads to $\sum_{n=1}^{K}\Sigma_{mn}=1,~\forall m$.
@@
The policy population gradually increases over time. When policy $m$ is selected for training, only policy $n$ $(n\leq m)$ has already been trained and holds meaningful outcomes. Other policies are inactive policies. As a result, we exclusively select policy $n$, where $n\leq m$, to serve as the opponent for policy $m$. ∎
3.2.2. Vanilla Self-Play
@@
In vanilla self-play (Samuel, 1959), agents are trained by competing against their latest versions. Thus, the MSS of vanilla self-play:
@@
(5) $\textit{MSS}(\mathcal{P})_{mn}=\left\{\begin{array}[]{l}1,~\text{if}~n=m-1\\ 0,~\text{otherwise}\end{array}\right.$.
@@
Regardless of $\mathcal{P}$, the MSS produces the same interaction matrix. Although this MSS is straightforward, it is utilized in many other algorithms, so we refer to it as vanilla MSS. Vanilla self-play is effective in transitive games, but it can lead the agent to cyclic learning patterns in non-transitive games. It may also lead to a nonstationary problem, causing the system to get stuck in a local optimum.
@@
A more modern version of vanilla self-play is to combine it with off-policy RL. In each iteration, the policy to be trained, $\pi_{[k]}^{h}$, is initialized using the previous policy $\pi_{[k-1]}(\cdot|h(k-1))$. Typically, each player’s policy is fixed, and samples are collected into a buffer by vanilla self-play. These samples remain useful for training, even if they were generated by an earlier policy. This allows the agent to leverage past experiences and maintain more stationary training. Once the policy is updated, it will be added to the policy population.
Moreover, vanilla self-play is frequently used in conjunction with Monte Carlo Tree Search (MCTS), and we provide a detailed discussion in Sec. 4.1.1, using the game of Go as a representative application.
3.2.3. Fictitious Play
@@
Fictitious Play (FP) (Brown, 1951) is a learning algorithm in game theory where each player best responds to the empirical frequency of the strategies used by the opponent. If the opponent’s strategy is static, FP can find the NE. Based on FP intuition, Fictitious Self-Play (FSP) (Heinrich et al., 2015) is introduced to make agents play against past versions of themselves to learn optimal strategies to improve the robustness of vanilla self-play. Neural Fictitious Self-Play (NFSP) (Heinrich and Silver, 2016) is a modern variant that combines FSP with deep learning techniques. It uses neural networks to approximate the BRs. Both NFSP and FSP utilize two distinct types of memory: one to store the agent’s own behavioral data, and the other to record observed opponent behavior.
@@
More recent approaches (Lanctot et al., 2017) adopt FP in meta-games, where the opponent’s average strategy is approximated directly, eliminating the need to maintain two separate types of memory. In this setting, the MSS of FP is defined as:
@@
(6) $\textit{MSS}(\mathcal{P})_{mn}=\left\{\begin{array}[]{l}\frac{1}{m-1},~\text{if}~n\leq m-1\\ 0,~\text{otherwise}\end{array}\right.$.
@@
In FP, the MSS continues to generate a constant interaction matrix. Compared to vanilla self-play, this approach enhances the robustness of the policy by sampling older versions of its own policies to be used as opponent strategies.
3.2.4. $\delta$-Uniform Self-Play
$\delta$-uniform self-play, introduced by Bansal et al. (2018), uses the hyper-parameter $\delta$, ranging from 0 to 1, to select the most recent $1-\delta$ percentage of policies for uniform sampling to generate opponent policies. When policy $m$ is in the training phase, according to Corollary 3.1, only the previous m1m-1 policies are relevant. To select opponents for policy $m$, we sample from the range of policies between $\lceil\delta(m-1)\rceil,m-1]$, where $\lceil\cdot\rceil$ denotes the ceiling function, which rounds up the input to the nearest integer greater than or equal to it. When $\delta=0$, the system retains the complete historical memory, whereas $\delta=1$ implies that only the most recent policy is utilized. Thus, the MSS of $\delta$-uniform self-play:
@@ 3.2.4. $\delta$-Uniform Self-Play
(7) $\textit{MSS}(\mathcal{P})_{mn}=\left\{\begin{array}[]{l}\frac{1}{f(m)},~\text{if}~\lceil\delta(m-1)\rceil\leq n\leq m-1\\ 0,~\text{otherwise}\end{array}\right$.,
(8) $f(m)=m-\lceil\delta(m-1)\rceil.$
@@
In $\delta$-uniform self-play, The MSS generates a constant interaction matrix. Specially, if $\delta=0$, it corresponds to FP, and if $\delta=1$, it corresponds to vanilla self-play.
3.2.5. Prioritized Fictitious Self-Play
@@
Prioritized Fictitious Self-Play (PFSP)  (Vinyals et al., 2019) leverages a priority function to allocate a higher probability of selection to agents with higher priorities. Here, $\mathcal{P}$ represents the winning rates, specifically defined as $\mathcal{P}_{mn}=\text{Prob}(\pi_{[m]}~\text{beats}~\pi_{[n]})$. The MSS of PFSP is given by Algo. 5. The function $f:[0,1]\rightarrow[0,\infty)$ is a priority function. For example, $f(x)=(1-x)^{p}$ with $p>0$ indicates that stronger policies against the currently being trained policy have a higher chance of being chosen as opponents. Alternatively, f=x(1x)f=x(1-x) implies that players of similar levels are more likely to be chosen as opponents. Additionally, in broader terms, $\mathcal{P}$ can also be assessed by other metrics. A similar MSS in  Berner et al. (2019) can be utilized to allocate higher probabilities to strategies that perform better or are more challenging to defeat.
@@
2:  $\sigma_{m+1,n}\leftarrow\frac{f(\mathcal{P}_{m,n})}{\sum_{n\leq m}f(\mathcal{P}_{m,n})},~\text{if}~n\leq m$
3:  Append zeros to $\sigma_{m+1}$ until its length is K.
@@ 3.3. PSRO Series
3.3.1. Integration into Our Framework
The PSRO series of algorithms can also be integrated into our proposed framework (Algo. 1). First, similar to traditional self-play algorithms, we also utilize lazy initialization to initialize $\Pi$. Second, we also set $E=1$ and $K$ can be considered as the upper limit for the policy population size in the original PSRO algorithms. Third, in the context of the PSRO series of algorithms, the strategy of our player $\pi_{[k]}^{h}$ can also be initialized in a general manner. Fourth, we set $h(k)=\emptyset$ since the PSRO series of algorithms do not use any policy condition function for their policies. Fifth, it’s crucial to highlight that our framework diverges from the traditional PSRO model (Lanctot et al., 2017) in how $\sigma_{[k]}$ is defined. In contrast to being the meta-strategy for policy, in our framework, $\sigma_{[k]}$ is the opponent sampling strategy. It means that $\sigma_{[k]}$ here represents the opponent’s meta-strategy against policy $k$ for the PSRO series of algorithms. Sixth, compared with traditional self-play methods, the MSSs of the PSRO series are often more complex. For example, some MSSs incorporate concepts from different types of game equilibria (McMahan et al., 2003; Muller et al., 2020; Marris et al., 2021). Lastly, we also simply follow Assumption 1. We can derive the Corollary 3.2 using a similar proof as Corollary 3.1.
@@ 3.3.1. Integration into Our Framework
3.3.2. Double Oracle
Double Oracle (DO) (McMahan et al., 2003) can be only applied to two-player normal-form games. In this context, we can utilize the payoff matrix as the evaluation matrix. The interaction matrix can be initialized with all zeros, reflecting the initial absence of interactions between strategies. The MSS of DO can then be outlined as described in Algo. 6. The opponent sampling strategy $\sigma_{[k]}$ corresponds to the opponent’s NE strategy of the restricted game. Therefore, the oracle in DO is a BR rather than an ABR, computing the BR against the current NE opponent strategy of the restricted game. In two-player normal-form settings, DO theoretically can achieve the NE of the full game.
@@ 3.3.2. Double Oracle
[…]
3:  Append zeros to $\sigma_{[k+1]}$ until its length is K.
@@
We take rock-paper-scissors as an example, a symmetric two-player game. The policy condition function $h(k)=\emptyset$. The policy population size $K=3$. The policy population $\Pi$ uses lazy initialization, and the interaction matrix $\Sigma$ is initialized with all zeros. Assume that in the first iteration, the policy $\pi_{1}$ is set to rock. Obviously, the NE-based MSS returns $\sigma_{2}=(1,0,0)$. In the second iteration, the BR oracle adds paper to $\Pi$ as $\pi_{2}$. Consequently, the NE-based MSS yields $\sigma_{3}=(0,1,0)$, since paper always beats rock. In the third iteration, noting that $\sigma_{3}$ selects only paper and never rock, the BR oracle adds scissors as $\pi_{3}$. The final interaction matrix is:
(9) $\Sigma=\begin{bmatrix}0&0&0\\ 1&0&0\\ 0&1&0\\ \end{bmatrix}$
@@
Finally, the NE over this policy population becomes the uniform distribution over rock, paper, and scissors, indicating convergence. This process demonstrates how the DO algorithm iteratively expands the strategy set within our framework. Also, using DO as an example, we show that, compared with traditional self-play algorithms, the PSRO series employs more sophisticated MSSs to better capture the essence of the game, albeit at a higher computational cost.
3.3.3. PSRO
@@
PSRO (Lanctot et al., 2017) extends DO to more complex games beyond just two-player normal-form games. This approach first introduces the concept of the MSS, which is a broader concept than simply computing NE. The MSS framework allows for a more flexible representation of strategic solutions in various game settings. Many variants of PSRO focus on designing new MSSs to better capture different aspects of strategic play in these more complex games. In the original PSRO framework, the oracle is computed using RL techniques, similar to those described in Algorithm 2. This allows the algorithm to effectively handle large and intricate strategy spaces, making it applicable to various game scenarios.
3.3.4. PSRO Variants
@@
The traditional PSRO algorithm has inspired numerous extensions in recent research. Although the original PSRO aligns well with our proposed framework, its variants may fall outside its scope; however, they remain noteworthy and merit discussion. The first line of PSRO variants focuses on accelerating convergence. By establishing a hierarchical structure of RL workers, Pipeline PSRO (McAleer et al., 2020) achieves the parallelization of PSRO and simultaneously provides guarantees of convergence. Efficient PSRO (Zhou et al., 2022) introduces the unrestricted-restricted (URR) game to narrow the selection of opponent policies, thereby preventing the need for meta-game simulations in traditional PSRO. Similar to Pipeline PSRO, Efficient PSRO is equipped to solve URR in parallel. In addition, unlike traditional PSRO, which confines the integration of population strategies to the game’s initial state, Extensive-Form Double Oracle (XDO) (McAleer et al., 2021) allows this integration across all information states. It ensures a linear convergence to an approximate NE based on the number of information states, enhancing its tractability in extensive-form games. Online Double Oracle (ODO) (Dinh et al., 2022) integrates the no-regret analysis of online learning with DO to improve both the convergence rate to an NE and the average payoff. Anytime PSRO (McAleer et al., 2022b) and Self-play PSRO (McAleer et al., 2022a) are designed to incorporate less exploitable policies into the policy population, facilitating faster convergence.
In parallel, other works aim to make PSRO more computationally tractable in different game settings. As the number of agents grows, determining BRs becomes exponentially challenging. Mean-field PSRO (Muller et al., 2022) has been introduced to address this complexity in mean-field games. In addition, due to the computational intractability of solving NE in multi-player general-sum games (Daskalakis et al., 2009; Daskalakis, 2013) and the selection problem in NE (Goldberg et al., 2013)Muller et al. (2020) put forward $\alpha$-PSRO. Instead of NE, $\alpha$-rank (Omidshafiei et al., 2019), which is unique and polynomial-time solvable in multi-player general-sum games, is introduced in the MSS, and a preference-based best response (PBR) oracle is incorporated in the approach. Similar to $\alpha$-PSRO, Marris et al. (2021) proposes the Joint Policy-Space Response Oracle (JPSRO) to tackle multi-player general-sum games. It utilizes CE and coarse correlated equilibrium (CCE) as alternatives to NE to make it computationally tractable in multi-player general-sum games and designs the CE and CCE-based MSS.
@@
Other lines of research focus on promoting policy diversity, especially in non-transitive games where a single policy is often insufficient. Balduzzi et al. (2019) introduces an open-ended framework in two-player zero-sum games. This framework enhances the diversity of the strategy population and introduces a gamescape, which geometrically represents the latent objectives in games for open-ended learning. The study proposed two algorithms: Nash response $\text{PSRO}_{\text{N}}$ and rectified Nash response $\text{PSRO}_{\text{rN}}$. Both algorithms utilize an asymmetric payoff matrix as their performance evaluation metric. Similarly to DO, they employ the Nash-based MSS (Algo. 6). Compared to $\text{PSRO}_{\text{N}}$, $\text{PSRO}_{\text{rN}}$ incorporates an additional step within the ABR oracle to focus on the opponents they beat or tie and ignore the opponents they lose to. Perez-Nieves et al. (2021) uses determinantal point process (Kulesza et al., 2012) to assess diversity and introduces diverse PSRO by incorporating a diversity term into the PSRO oracle. This modification can also be easily implemented in FP and $\alpha$-PSRO. Similarly,  Liu et al. (2021) introduces the concepts of behavioral diversity and response diversity and incorporates them into the PSRO oracle. Policy Space Diversity PSRO (Yao et al., 2023) defines a diversity metric named population exploitability that helps to achieve a full-game NE.
3.3.5. R-NaD